OpenStudy (sburchette):

I have a proof of a theorem for abstract algebra, but I want to ensure that it is valid. The theorem states that gcd(a,bc)=1 iff gcd(a,b)=1 and gcd(a,c)=1. My proof goes as follows: Assume that gcd(a, bc)=1 for some integers a, b, and c. Then, there must exist integers s and t such that as+bct =1. Since 1 can be written as a linear combination of a and b, gcd(a,b)=1. 1 can also be written as a linear combination of a and c, so gcd(a,c)=1. Now assume gcd(a,b)=1 and gcd(a,c)=1. There must exist integers s, t, u, and v such that 1=as+bt and 1=au+cv. Multiplying both equations, we have 1*1 = (as+bt)(au+cv). From this, we have 1=a^2su+acv+btau+btcv. Factoring out an a from the first three terms, we have that 1=a(asu+cv+btu)+bc(tv). Since 1 has been written has a linear combination of a and bc, we have gcd(a,bc)=1. QED.

OpenStudy (loser66):

post it, I tag someone, when they came online, they can check @oldrin.bataku @pgpilot326 @Wallach

OpenStudy (dan815):

how can you say that is a linear combination, isnt it now dependant on a and b

OpenStudy (loser66):

no, dan, it's linear. because gcd(a,bc) =1

OpenStudy (anonymous):

\(asu+cv+btu=m\in \mathbb{Z}\text{ and } tv=n\in\mathbb{Z}\) proof looks good to me

OpenStudy (anonymous):

why medal me? i did'n do the proof. did'n... ha ha. i mean didn't.

OpenStudy (anonymous):

aw

OpenStudy (dan815):

give it to him

OpenStudy (sburchette):

oops, didn't mean to hit undo

OpenStudy (anonymous):

por que?

OpenStudy (anonymous):

i mean why the medals?

OpenStudy (anonymous):

give em to @Loser66 it's his proof.

OpenStudy (loser66):

hahahah, not me, I didn't start abstract algebra yet, next week, not now

OpenStudy (anonymous):

i haven't taken it either...

OpenStudy (dan815):

if gcd (a,bc)=1 if gcd(a,c)=1 and gcd (a,b)=1 then (a)+(bc) = n(a1+b1c2) suppose n not equal to 1 then gcd(a,c) .. a+c = n(a1+bc2), therefore gcd(a,c) not equal =1 similarly gcd(a,b) proof by contradiction , therefore if gcd(a,b)=1 and gcd(a,c)=1, then gcd(a,bc)=1

OpenStudy (dan815):

where a1,b1,and c2 E Z